### v2

# Entropic additive energy and entropy inequalities for sums and products

Abstract
Following a growing number of studies that, over the past 15 years, have established entropy inequalities via ideas and tools from additive combinatorics, in this work we obtain a number of new bounds for the differential entropy of sums, products, and sum-product combinations of continuous random variables. Partly motivated by recent work by Goh on the discrete entropic version of the notion of “additive energy”, we introduce the additive energy of pairs of continuous random variables and prove various versions of the statement that “the additive energy is large if and only if the entropy of the sum is small”, along with a version of the Balog–Szemerédi–Gowers theorem for differential entropy. Then, motivated in part by recent work by Máthé and O’Regan, we establish a series of new differential entropy inequalities for products and sum-product combinations of continuous random variables. In particular, we prove a new, general, ring Plünnecke–Ruzsa entropy inequality. We briefly return to the case of discrete entropy and provide a characterization of discrete random variables with “large doubling”, analogous to Tao’s Freiman-type inverse sumset theory for the case of small doubling. Finally, we consider the natural entropic analog of the Erdős–Szemerédi sum-product phenomenon for integer-valued random variables. We show that, if it does hold, then the range of parameters for which it does would necessarily be significantly more restricted than its anticipated combinatorial counterpart.
1 Introduction
1.1 Additive combinatorics and entropy inequalities
The field of additive combinatorics [26] studies occurrences of additive structures, such as arithmetic progressions, in subsets of abelian groups like the additive group of the integers $({\mathbb{Z}},+)$. The Plünnecke–Ruzsa sumset theory is an important sub-field of additive combinatorics that studies sumset inequalities, which are used in many branches of combinatorics as well as in other areas of mathematics.
The sumset $A+B$ of two discrete subsets $A$ and $B$ of an additive group $G$, i.e., an abelian group $(G,+)$, is $A+B=\left\{a+b:a\in A,b\in B\right\}$. Sumset inequalities include the trivial bound
$\left\lvert A+B\right\rvert\geq\max\left\{\left\lvert A\right\rvert,\left\lvert B\right\rvert\right\}$, (1)
as well as many more nontrivial and subtle results. For example, if we similarly define the difference set $A-B=\left\{a-b:a\in A,b\in B\right\}$, then the Ruzsa triangle inequality [20] says that the cardinality of ACA-C can be bounded in terms of the cardinalities of ABA-B and BCB-C via
$\left\lvert A-C\right\rvert\leq\frac{\left\lvert A-B\right\rvert\left\lvert B-C\right\rvert}{\left\lvert B\right\rvert}$, (2)
and the sum-difference inequality [22] compares the cardinality of $A+B$ with that of ABA-B:
$\left\lvert A+B\right\rvert\leq\frac{\left\lvert A-B\right\rvert^{3}}{\left\lvert A\right\rvert\left\lvert B\right\rvert}$. (3)
Shannon’s Asymptotic Equipartition Property (AEP) motivates an interesting connection between sumset inequalities and information theory. Let $X$ be a discrete random variable with values in a set $A$, and with entropy $H=H(X)$, in nats. Suppose $X_{1},X_{2},\ldots$ are independent and identically distributed (i.i.d.) copies of $X$. Then the AEP tells us that there is a subset $B^{*}_{n}$ of $A^{n}$ such that, for large $n$, $B_{n}^{*}$ carries essentially all of the probability of the distribution of $(X_{1},\ldots,X_{n})$, i.e., ${\mathbb{P}}((X_{1},\ldots,X_{n})\in B^{*}_{n})=1-o(1)$, and its cardinality is minimal: It contains $|B^{*}_{n}|=e^{n(H+o(1))}$ elements, while any other subset $B_{n}$ of $A^{n}$ with ${\mathbb{P}}((X_{1},\ldots,X_{n})\in B_{n})=1-o(1)$ is necessarily at least as large, $|B_{n}|\geq e^{n(H-o(1))}.$ Therefore, the random variables $(X_{1},\ldots,X_{n})$ are essentially supported on a set of size $\doteq e^{nH}$, which is typically much smaller than the size $|A|^{n}=e^{n\log|A|}$ of all of $A^{n}$. In this sense, we may think of $e^{H}$ as the cardinality of the “essential support” of $X$, and consequently we can interpret the entropy $H(X)$ of a discrete random variable $X$ as a “probabilistic analog” of the log-cardinality of a discrete set $A$.
Thus motivated, Ruzsa [23] in 2009 and Tao [25] in 2010 began a systematic exploration of this entropy/cardinality correspondence, in the context of sumset inequalities. Specifically, they asked whether, starting with an arbitrary sumset bound and replacing discrete subsets of an additive group with independent discrete random variables, and log-cardinalities by entropies, would lead to a legitimate new entropy inequality. Although far from obvious, the answer generally turns out to be “yes”. For example, for independent random variables $X$ and $Y$, the simple bound in (1) corresponds to the familiar, elementary entropy inequality
$H(X+Y)\geq\max\left\{H(X),H(Y)\right\}.$
Similarly [25], for independent $X,Y,Z$, the entropic version of the Ruzsa triangle inequality in (2) is
$H(X-Z)+H(Y)\leq H(X-Y)+H(Y-Z),$
and the sum-difference inequality in (3) becomes
$H(X+Y)+H(X)+H(Y)\leq 3H(X-Y).$
This correspondence is rich and nontrivial, with neither the combinatorial nor the entropic version of these bounds being formally stronger than the other.
The entropy inequalities resulting from the entropy/cardinality correspondence have been extended and generalized along several directions, e.g., in [16, 17, 14, 4, 5, 9], and they have found numerous applications in core information-theoretic problems; see, e.g., [13, 24, 10, 15, 27]. In the reverse direction, entropic sumset bounds have also been finding important applications to questions in combinatorics. For example, the recent resolution by Gowers, Green, Manners and Tao [8, 7] of two important cases of the polynomial Freiman–Ruzsa conjecture critically depends on the entropic formulation of this purely combinatorial problem.
1.2 Differential entropy inequalities
Given the wealth and utility of the discrete entropy inequalities produced via the entropy/cardinality correspondence, it is natural to ask whether they extend to the case of the differential entropy for continuous random variables, i.e., for real-valued random variables with densities with respect to Lebesgue measure on $\mathbb{R}$}. Additional motivation for this comes from the continuous version of the AEP: Similarly to the discrete case, the differential entropy $h=h(X)$ of a continuous random variable $X$ may be interpreted as the size (in the sense of Lebesgue measure) of the essential support of $X$. As it turns out, it is generally possible to translate discrete entropic sumset inequalities to corresponding differential entropy bounds, but it is not entirely straightforward.
An essential element in the proofs of the discrete entropic versions of most sumset inequalities is the following observation: Consider two arbitrary discrete random variables $X,Y$. Suppose that the random variable $Z$ is a function of the pair $(X,Y)$, i.e., $Z=\phi(X,Y)$ for some function $\phi$, and that the random variable $W$ can be expressed as a function of either $X$ or $Y$, i.e., $W=f(X)=g(Y)$ for appropriate functions $f,g$. Then it is not hard to show [25] that
$H(W)+H(Z)\leq H(X)+H(Y).$
This property is sometimes referred to as the functional submodularity of Shannon entropy.
It is easy to come up with examples showing that functional submodularity fails in the case of differential entropy [12]. Moreover, the steps in which functional submodularity is used in the original proofs of discrete entropic sumset bounds involve intermediate inequalities that are not valid for continuous random variables and differential entropy. Nevertheless, as was shown in [12] and [16], essentially all of the earlier discrete entropic results obtained by Ruzsa [23], Tao [25], and others have near-identical continuous analogs, but instead of functional submodularity, the key ingredient in their proofs is the data processing property of mutual information.
Indeed, for any independent continuous random variables $X,Y,Z$, we have the elementary bound
$h(X+Y)\geq\max\{h(X),h(Y)\},$
the Ruzsa triangle inequality [12] becomes
$h(X-Z)+h(Y)\leq h(X-Y)+h(Y-Z),$
and the continuous analog of the sum-difference inequality [12] is simply
$h(X+Y)+h(X)+h(Y)\leq 3h(X+Y).$
In this paper, we continue the exploration of continuous versions of discrete entropy sumset bounds, largely motivated by recent results obtained by Goh [5] and by Máthé and O’Regan [18]. In the process, we also examine some questions regarding the entropy of sums and products of discrete random variables. As in the earlier work [12, 16], the main obstacle to obtaining continuous generalizations of the discrete bounds in [5] and [18] is the fact that their proofs rely heavily on the functional submodularity property of discrete entropy. Therefore, substantially new arguments are required, and the data processing property of mutual information again plays a key role.
1.3 Outline of main results
After a preliminary discussion of earlier work in Section 2, our main results, briefly outlined below, are developed in Sections 38.
Additive energy and the differential entropy of sums. In Section 3 we define the additive energy, $a(X,Y)$, for pairs of continuous random variables $(X,Y)$, and we observe that it can be equivalently expressed as a(X,Y)=2h(X,Y)h(X+Y)a(X,Y)=2h(X,Y)-h(X+Y). This representation immediately leads to the main subject of this section, which is to establish various quantitative versions of the statement that “if $a(X,Y)$ is large (close to its upper bound) then $h(X+Y)$ is small (close to its lower bound)”, as well as its converse. In this sense, the additive energy $a(X,Y)$ can naturally be interpreted as a measure of the “additive structure” present in $X$ and $Y$, in close analogy with its discrete counterpart examined in [5].
Balog–Szemerédi–Gowers theorem for differential entropy. In Section 4, we prove a version of the Balog–Szemerédi–Gowers (BSG) theorem for differential entropy and discuss its relation with earlier (both discrete and continuous) versions. Roughly speaking, the BSG theorem is the important assertion that, if the additive energy $a(X,Y)$ is appropriately “large”, then we can find a conditioning random variable $Z$ such that $X$ and $Y$ are approximately conditionally independent given $Z$, the values of the entropies $h(X|Z)$ and $h(Y|Z)$ are close to those of the original entropies $h(X)$ and $h(Y)$, respectively, and $h(X+Y|Z)$ is appropriately small.
Stability of large discrete entropic doubling. In Section 5 we briefly return to the discrete setting, and we consider the problem of characterizing discrete random variables $X$ with the property that, when $X$ and $X^{\prime}$ are i.i.d., then the entropy $H(X+X^{\prime})$ of their sum is close to its maximum possible value, $2H(X)$. This complements the Freiman-type inverse sumset theory of Tao for the case when $H(X+X^{\prime})$ is close to its obvious lower bound, $H(X)$. Interestingly, we find that, when $H(X+X^{\prime})$ is appropriately large, then $X$ is necessarily approximately supported on a Sidon set, i.e., a set $A$ for which the sums $a+b$ of all pairs of elements $a,b\in A$ are unique.
Differential entropy bounds for random products. In Section 6, we discuss how many of the entropic sumset inequalities in earlier sections and in earlier work can be generalized for products rather than sums of (both discrete and continuous) random variables. The results in this section mostly follow directly from the general development in [16], and they provide useful tools for the new inequalities developed next.
Sum-product inequalities for differential entropy. The previous section provides important tools for Section 7, which extends recent work of Máthé and O’Regan [18] on entropic sum-product entropic inequalities, namely, entropy bounds that involve both sums and products of random variables. The main result in this section is a new and general “ring Plünnecke–Ruzsa entropy inequality”. It gives an upper bound on the entropy of sums of products of i.i.d. random variables, in terms of their marginal entropy and their associated pairwise “doubling” constants, introduced and discussed in Section 6.2.
On the entropic Erdős–Szemerédi sum-product phenomenon. Finally, in Section 8 we ask whether an entropic Erdős–Szemerédi sum-product phenomenon exists for integer-valued random variables. Specifically, we ask whether it is necessarily the case that, if $X$ and $X^{\prime}$ are i.i.d. with values in $\mathbb{Z}$}, then at least one of $H(X+X^{\prime})$ and $H(XX^{\prime})$ must be significantly larger than $H(X)$. We do not present any positive results, but we show, via example, that the range of the validity of the naive version of this conjecture must necessarily be more restricted than both what is known and what is conjectured to hold in the combinatorial case.
2 Preliminaries and some prior work
2.1 Combinatorial background
In this section we describe the main concepts and results from additive combinatorics that motivate our subsequent entropic bounds. In particular, we introduce the combinatorial versions of the Ruzsa distance and the additive energy, and we describe those of their properties that will be relevant to the development of our results in Sections 38. All of the results in this section can be found, e.g., in the text [26].
Let $G$ be an additive group, that is, an arbitrary abelian group $(G,+)$. The Ruzsa distance $d(A,B)$ between two finite subsets $A$ and $B$ of $G$ is
$d(A,B)=\log\frac{\left\lvert A-B\right\rvert}{\sqrt{\left\lvert A\right\rvert\left\lvert B\right\rvert}}$, (4)
where, throughout the paper, $\log$’ denotes the natural logarithm. Note that the Ruzsa distance is not a metric, as $d(A,A)\neq 0$ in general. In this notation, the Ruzsa triangle inequality in (2) becomes
$d(A,C)\leq d(A,B)+d(B,C).$ (5)
Similarly, the sum-difference inequality in (3) can be expressed as
$d(A,-B)\leq 3d(A,B).$ (6)
In the case $A=B$, the following tighter bounds are available, sometimes referred to as the doubling-difference inequality
$\frac{1}{2}d(A,A)\leq d(A,-A)\leq 2d(A,A).$ (7)
The doubling constant of a finite set $A\subset G$ is simply the ratio:
$s(A)=\frac{|A+A|}{|A|}.$ (8)
More generally, for the $k$-fold addition of a set $A$ to itself, we write
$kA=\left\{a_{1}+a_{2}+\cdots+a_{k}:a_{1},a_{2},\dots,a_{k}\in A\right\},$
for any nonnegative integer $k$. The Plünnecke–Ruzsa inequality [19, 21] is a fundamental result in additive combinatorics, which states the following. For any finite subsets $A,B$ of $G$, if $\left\lvert A+B\right\rvert\leq K\left\lvert A\right\rvert$ for some $K>0$, then for all nonnegative integers $m,n$ we have:
$\left\lvert nB-mB\right\rvert\leq K^{n+m}\left\lvert A\right\rvert$. (9)
This bound is often used when $A=B$, where the best constant $K$ is the doubling constant $s(A)$.
The additive energy of two subsets $A,B$ of $G$ is
$E(A,B)=\left\lvert\{(a_{1},a_{2},b_{1},b_{2})\in A\times A\times B\times B:a_{1}+b_{1}=a_{2}+b_{2})\}\right\rvert$. (10)
Although it is not an “energy” function in the sense of physics, the natural interpretation of $E(A,B)$ is as a measure of the degree of additive structure present in $A+B$; see, e.g., [26] for extensive discussion of its properties. For example, $E(A,A)$ is always bounded above by $|A|^{3}$, and it is close to $|A|^{3}$ if and only if $A$ is close to an arithmetic progression.
An elementary but important inequality for $E(A,B)$ is:
$E(A,B)\geq\frac{\left\lvert A\right\rvert^{2}\left\lvert B\right\rvert^{2}}{\left\lvert A+B\right\rvert}$. (11)
One application of this inequality is in showing that a small sumset $A+B$ implies large additive energy $E(A,B)$, further reinforcing the interpretation $E(A,B)$ described above. To see this, note that since $\left\lvert A+B\right\rvert\geq\max\left\{\left\lvert A\right\rvert,\left\lvert B\right\rvert\right\}$, we always have $\left\lvert A+B\right\rvert\geq\left\lvert A\right\rvert^{1/2}\left\lvert B\right\rvert^{1/2}$. We can thus consider $\left\lvert A+B\right\rvert$ to be “small” if $\left\lvert A+B\right\rvert\leq C\left\lvert A\right\rvert^{1/2}\left\lvert B\right\rvert^{1/2}$ for some constant $C>1$. Then (11) leads to the implication:
$\left\lvert A+B\right\rvert\leq C\left\lvert A\right\rvert^{1/2}\left\lvert B\right\rvert^{1/2}\quad\Rightarrow\quad E(A,B)\geq\frac{1}{C}\left\lvert A\right\rvert^{3/2}\left\lvert B\right\rvert^{3/2}.$ (12)
2.2 Discrete entropy, Ruzsa distance and additive energy
Here we describe the entropic analogs of the Ruzsa distance and the additive energy, together with some of the main properties, along the lines of the entropy/cardinality correspondence. The results in this section can be found in [25] and [5]. The entropy of an arbitrary discrete random variable $X$ with probability mass function $P$ on a discrete alphabet $A$ is defined, as usual, as $H(X)={\mathbb{E}}[-\log P(X)]$. In all our results throughout the paper, it is implicitly assumed that all entropies that appear in the statements are finite.
Let $X,Y$ be two discrete random variables with values in an additive group $G$. In analogy with (4), the entropic Ruzsa distance $d(X,Y)$ between $X$ and $Y$ is defined as
$d(X,Y)=H(X^{\prime}-Y^{\prime})-\frac{1}{2}H(X)-\frac{1}{2}H(Y),$ (13)
where $X^{\prime}$ and $Y^{\prime}$ are independent copies of $X$ and $Y$, respectively. Once again, we note that this is not a metric, as $d(X,X)\neq 0$ in general. Also we observe that $d(X,Y)$ does not depend on the joint distribution of $(X,Y)$, only on their marginals.
In this notation, the entropic version of the Ruzsa triangle inequality (5) states that, for any three discrete random variables $X,Y,Z$,
$d(X,Z)\leq d(X,Y)+d(Y,Z),$ (14)
the entropic version of the sum-difference inequality (6) is
$d(X,-Y)\leq 3d(X,Y),$ (15)
and the entropic doubling-difference inequality corresponding to (7) is
$\frac{1}{2}d(X,X)\leq d(X,-X)\leq 2d(X,X).$ (16)
An entropic analog of the Plünnecke–Ruzsa inequality (9) was proved by Tao [25]. It says that, if $X_{1},X_{2},\dots,X_{n},X_{1}^{\prime},X_{2}^{\prime},\dots,X_{m}^{\prime}$ are independent copies of a discrete random variable $X$ taking values in an additive group, and $H(X_{1}+X_{1}^{\prime})\leq H(X)+\log K$ for some $K\geq 1$, then,
$H(X_{1}+X_{2}+\cdots+X_{n}-X_{1}^{\prime}-X_{2}^{\prime}-\cdots-X_{m}^{\prime})\leq H(X)+O(n+m)\log K.$ (17)
This is formally weaker than (9) because it loses an absolute constant with respect to the exponent $n+m$, and also because the constant $K$ in the bound requires $X$ and $X^{\prime}$ to have the same distribution.
Continuing in the spirit of the entropy/cardinality correspondence, Goh [5] recently introduced a version of the additive energy of pairs of discrete random variables, and explored its properties. In analogy with the combinatorial version of additive energy (10), the entropic additive energy of two jointly distributed discrete random variables $X$ and $Y$ taking values in the same additive group is
$A(X,Y)=H(X_{1},Y_{1},X_{2},Y_{2},X+Y),$ (18)
where $(X_{1},Y_{1})$ and $(X_{2},Y_{2})$ are conditionally independent versions of $(X,Y)$ given $X+Y$. Intuitively, we can think of first generating $(X,Y)$ and then independently generating $(X_{1},Y_{1})$ and $(X_{2},Y_{2})$ from the same distribution as $(X,Y)$, conditioned on the event $\{X_{1}+Y_{1}=X_{2}+Y_{2}=X+Y\}$.
We will also find it convenient to use the following equivalent expression for $A(X,Y)$:
$A(X,Y)$ $=H(X_{1},Y_{1},X_{2},Y_{2},X+Y)$
$=H(X+Y)+H(X_{1},Y_{1}|X+Y)+H(X_{2},Y_{2}|X+Y)$
$=2H(X,Y|X+Y)+H(X+Y)$
$=2H(X,Y)-H(X+Y).$ (19)
This expression clearly leads to the same interpretation as for the combinatorial additive energy: Small entropy of the sum $X+Y$ implies large additive energy $A(X,Y)$. This interpretation is further reinforced by the entropic analog of the implication (12). Here we actually have that, if $X,Y$ are independent, then
$H(X+Y)\leq\frac{1}{2}H(X)+\frac{1}{2}H(Y)+\log C\quad\Leftrightarrow\quad A(X,Y)\geq\frac{3}{2}H(X)+\frac{3}{2}H(Y)-\log C.$ (20)
Moreover, for arbitrary pairs of (not necessarily independent) discrete random variables $(X,Y)$, the reverse implication remains valid: If the additive energy $A(X,Y)$ is large, then $H(X+Y)$ is small.
Both in the combinatorial case and in the entropy case, the terms “large” and “small” are understood in the sense that the relevant constants $C$ are fixed while the set cardinalities and the random variable entropies may be arbitrarily large.
2.3 Differential entropy and entropic Ruzsa distance
As was shown in [12], all of the entropy bounds of the previous section extend to continuous random variables. Apart from the proof of the sum-difference inequality which is new, all other results in this section can be found in [12]. The differential entropy $h(X)$ of a continuous random vector $X$ with density $f$ in ${\mathbb{R}}^{d}$, or, equivalently, of a random vector $X=(X_{1},\ldots,X_{d})$ in ${\mathbb{R}}^{d}$ with joint density $f$, is
$h(X)=-\int_{{\mathbb{R}}^{d}}f(x)\log f(x)\,dx={\mathbb{E}}[-\log f(X)],$
whenever the integral exists. Otherwise, we let $h(X)=$-$\infty$. Throughout the paper, the differential entropy of any continuous random variable or random vector appearing in the statement of any of our results is assumed to exist and be finite.
Similarly to the discrete case (13), the Ruzsa distance between two continuous random variables $X,Y$ is
$d(X,Y)=h(X^{\prime}-Y^{\prime})-\frac{1}{2}h(X)-\frac{1}{2}h(Y),$ (21)
where $X^{\prime}$ and $Y^{\prime}$ are independent copies of $X$ and $Y$, respectively. Then, for arbitrary continuous random variables $X,Y,Z$, the Ruzsa triangle inequality (14), the sum-difference inequality (15), and the doubling-difference inequality (16) remain valid exactly as before. Another useful bound in the same spirit is the submodularity for sums inequality, which says that, if $X,Y,Z$ are independent random vectors, then
$h(X+Y+Z)+h(Y)\leq h(X+Y)+h(Y+Z).$ (22)
The obvious analog of (22) is also true in the case of discrete entropy [11].
The proof of the sum-difference inequality for differential entropy in [12] used a construction similar to that in Tao’s proof of the discrete version [25], with the steps that employed functional submodularity replaced by different intermediate inequalities that follow from the data processing property of mutual information. Although both these proofs are quite involved, we note that the submodularity-for-sums inequality can be used to provide a very short proof that works in both the discrete and continuous case.
Proof of the sum-difference inequality. We prove the result for the continuous case; the discrete case follows from an identical argument. Without loss of generality, assume $X$ and $Y$ are independent, and let $X^{\prime}$ be an independent copy of $X$. Then
$d(X,-Y)$ $=h(X+Y)-\frac{1}{2}h(X)-\frac{1}{2}h(Y)$
$\leq h(X+Y-X^{\prime})-\frac{1}{2}h(X)-\frac{1}{2}h(Y)$
$\leq h(X-X^{\prime})+h(Y-X^{\prime})-\frac{3}{2}h(X)-\frac{1}{2}h(Y)$
$=d(X,X)+d(X,Y)$
$\leq 3d(X,Y).$
The three inequalities follow from the fact that convolution increases differential entropy, the submodularity for sums bound, and the Ruzsa triangle inequality, respectively.   $\Box$
In the same work [12], the following version of the Plünnecke–Ruzsa inequality for differential entropy is established: Suppose $X,Y_{1},Y_{2},\dots,Y_{n}$ are independent continuous random variables. If there are constants $K_{1},K_{2},\dots,K_{n}\geq 1$ satisfying $h(X+Y_{i})\leq h(X)+\log K_{i}$ for each $i$, then
$h(X+Y_{1}+Y_{2}+\cdots+Y_{n})\leq h(X)+\log(K_{1}K_{2}\cdots K_{n}).$ (23)
Using the doubling-difference inequality, this implies the differential entropy analog of (17) with the explicit factor $n+2m$ in place of the $O(n+m)$ term. Once again, we note that the proof immediately translates to the discrete case, and the corresponding inequality holds for discrete random variables.
3 Additive energy and the differential entropy of sums
In this section we define the additive energy $a(X,Y)$ for continuous random variables $X,Y$, and we establish various statements that elaborate on the intuition that large additive energy $a(X,Y)$ corresponds to small entropy $h(X+Y)$ of the sum of $X$ and $Y$. These results generalize and strengthen the corresponding bounds for the entropy of discrete random variables obtained by Goh [5]. As the proofs of the discrete bounds in [5] rely heavily on the functional submodularity property of discrete entropy, we take a different approach to proving the corresponding results for differential entropy.
The gist of our approach is the observation that the combinatorial versions of all the results we consider in this section follow from the basic bound (11) described in the Introduction. Therefore, we begin by establishing its natural entropic analog in Lemma 3.2. From this, we then derive the continuous analogs of all the relevant results obtained in [5]. This approach has some important advantages, that indeed apply to almost all the results obtained in this paper:
  • The resulting proofs do not rely on functional submodularity, so they apply to both the continuous and discrete cases.
  • The bounds obtained are often stronger than their earlier discrete counterpart, and the slack in the inequalities can sometimes be precisely quantified.
In particular, although our results in this section are stated for continuous random variables, they all apply verbatim to discrete random variables as well.
3.1 Additive energy for continuous random variables
Recall that, for an arbitrary pair of discrete random variables $(X,Y)$ with values in an additive group $G$, we defined the entropic additive energy $A(X,Y)$ in (18) as $A(X,Y)=H(X_{1},Y_{1},X_{2},Y_{2},X+Y),$ where $(X_{1},Y_{1})$ and $(X_{2},Y_{2})$ are conditionally independent versions of $(X,Y)$, given $X+Y$. But $Y_{1}$ and $Y_{2}$ can be determined from $(X_{1},X_{2},X+Y)$ so, if $X,Y$ are continuous random variables, then $h(X_{1},Y_{1},X_{2},Y_{2},X+Y)=$-$\infty$. On the other hand, the same argument shows that $A(X,Y)$ is also always equal to $H(X_{1},X_{2},X+Y)$. Therefore, we define the differential entropic additive energy of two jointly distributed continuous random variables $X$ and $Y$ as
$a(X,Y)=h(X_{1},X_{2},X+Y),$
where $(X_{1},Y_{1})$ and $(X_{2},Y_{2})$ are conditionally independent versions of $(X,Y)$, given $X+Y$.
As with the expression for $A(X,Y)$ derived in (19) in the discrete case, we will find the following alternative representation $a(X,Y)$ useful and sometimes simpler to work with.
Lemma 3.1.
For any pair $(X,Y)$ of continuous random variables:
a(X,Y)=2h(X,Y)h(X+Y).a(X,Y)=2h(X,Y)-h(X+Y).
Proof. Let $(X_{1},Y_{1})$ and $(X_{2},Y_{2})$ be conditionally independent versions of $(X,Y)$ given $X+Y$. By the chain rule and conditional independence, we have,
$h(X_{1},X_{2},X+Y)$ $=h(X_{1}|X+Y)+h(X_{2}|X+Y)+h(X+Y)$
$=h(X_{1},X+Y)-h(X+Y)+h(X_{2},X+Y)-h(X+Y)+h(X+Y)$
$=h(X_{1},Y_{1})+h(X_{2},Y_{2})-h(X+Y)$
$=2h(X,Y)-h(X+Y),$
as claimed.   $\Box$
As mentioned above, the starting point for all results in this section will be the following entropic analog of (11). Recall that the mutual information $I(X;Y)$ between two continuous random variables $X$ and $Y$ satisfies I(X;Y)=h(X)+h(Y)h(X,Y)I(X;Y)=h(X)+h(Y)-h(X,Y), and similarly in the discrete case.
Lemma 3.2.
Let $X$ and $Y$ be continuous random variables. Then
$h(X+Y)+a(X,Y)=2h(X,Y)\leq 2h(X)+2h(Y),$ (24)
where the slackness in the inequality is exactly $2I(X;Y)\geq 0$.
Note that the inequality in (24) appears to be in the reverse direction compared to its combinatorial counterpart (11). But when $X$ and $Y$ are independent – as dictated by the entropy/cardinality correspondence – we actually have equality in (24), which is of course consistent with (11).
3.2 Large entropic additive energy
In order to quantify what “large” means for additive energy, note that
$a(X,Y)=2h(X,Y)-h(X+Y)\leq 2h(X,Y)-h(X|Y)=h(X,Y)+h(Y),$
so, by symmetry,
$a(X,Y)\leq h(X,Y)+\min\{h(X),h(Y)\}\leq h(X)+h(Y)+\min\{h(X),h(Y)\}.$ (25)
Therefore, we can consider $a(X,Y)$ to be large if
$a(X,Y)\geq\frac{3}{2}h(X)+\frac{3}{2}h(Y)-\log C.$
Similarly, since
$h(X+Y)\geq\max\{h(X+Y|Y),h(X+Y|X)\}=\max\{h(X|Y),h(Y|X)\}\geq\frac{1}{2}h(X|Y)+\frac{1}{2}h(Y|X),$
we can consider $h(X+Y)$ to be small if
$h(X+Y)\leq\frac{1}{2}h(X|Y)+\frac{1}{2}h(Y|X)+\log C.$
The following two results show that large $a(X,Y)$ implies small $h(X+Y)$, with a partial converse if $X$ and $Y$ are only weakly dependent. Corollary 3.3 is the differential entropy version of the discrete entropy result in [5, Proposition 3].
Corollary 3.3.
Let $X,Y$ be continuous random variables. If
$a(X,Y)\geq\frac{3}{2}h(X)+\frac{3}{2}h(Y)-\log C,$ (26)
for some constant $C$, then
$h(X+Y)\leq\frac{1}{2}h(X)+\frac{1}{2}h(Y)+\log C.$ (27)
Conversely, if (27) holds and $X$ and $Y$ are weakly dependent in the sense that
$h(X,Y)\geq h(X)+h(Y)-\log C^{\prime},$ (28)
for some constant $C^{\prime}$, then
$a(X,Y)\geq\frac{3}{2}h(X)+\frac{3}{2}h(Y)-\log C-2\log C^{\prime}.$ (29)
In particular, if $X$ and $Y$ are independent, then (27) implies (26).
Proof. By Lemma 3.2, the bound (26) implies (27). Conversely, by  (27), (28), and Lemma 3.2, we have
$2h(X)+2h(Y)-2\log C^{\prime}$ $\leq 2h(X,Y)$
$=a(X,Y)+h(X+Y)$
$\leq a(X,Y)+\frac{1}{2}h(X)+\frac{1}{2}h(Y)+\log C,$
and rearranging yields (29).   $\Box$
If, instead of using Lemma 3.2, we use the mutual information $I(X;Y)$, we immediately obtain the following strengthening of Corollary 3.3 and of the corresponding discrete entropy result in [5, Proposition 3].
Corollary 3.4.
Let $X$ and $Y$ be continuous random variables and $C$ be a constant. Then
$a(X,Y)\geq\frac{3}{2}h(X)+\frac{3}{2}h(Y)-\log C,$
if and only if,
$h(X+Y)\leq\frac{1}{2}h(X)+\frac{1}{2}h(Y)+\log C-2I(X;Y).$
Proof. From the definitions we have
$a(X,Y)-\frac{3}{2}h(X)-\frac{3}{2}h(Y)$ $=2h(X,Y)-h(X+Y)-\frac{3}{2}h(X)-\frac{3}{2}h(Y)$
$=\frac{1}{2}h(X)+\frac{1}{2}h(Y)-2[h(X)+h(Y)-h(X,Y)]-h(X+Y)$
$=\frac{1}{2}h(X)+\frac{1}{2}h(Y)-2I(X;Y)-h(X+Y),$
which gives the claimed equivalence.   $\Box$
Next, we show that either of the statements in the above two corollaries – namely, $a(X,Y)$ being large in the sense of (26), or $h(X+Y)$ being small in the sense of (27) – is equivalent to the entropies $h(X)$ and $h(Y)$ being close in value. The following result generalizes and strengthens [5, Proposition 7] for discrete entropy.
Corollary 3.5.
Let $X$ and $Y$ be continuous random variables and $C$ be a constant. Then we have
$a(X,Y)\geq\frac{3}{2}h(X)+\frac{3}{2}h(Y)-\log C$ (30)
if and only if
$h(X)\leq h(Y)+2\log C-2I(X+Y;Y)+2I(X;Y),$ (31)
or, equivalently,
$h(Y)\leq h(X)+2\log C-2I(X+Y;X)+2I(X;Y).$ (32)
Under any, hence all, of these conditions,
$h(Y)-2\log C+2I(X+Y;X)-2I(X;Y)\leq h(X)\leq h(Y)+2\log C-2I(X+Y;Y)+2I(X;Y).$ (33)
In particular,
$\left\lvert h(X)-h(Y)\right\rvert\leq 2\log C+2I(X;Y)-2\min\{I(X+Y;X),I(X+Y;Y)\}.$ (34)
Proof. Using the bound (30) in the equality part of Lemma 3.2 yields
$\frac{3}{2}h(X)+\frac{3}{2}h(Y)-\log C+h(X+Y)\leq h(X)+h(Y)+h(X|Y)+h(Y|X).$
Since $h(X+Y)=h(X|Y)+I(X+Y;Y)$ and h(Y|X)=h(Y)I(X;Y)h(Y|X)=h(Y)-I(X;Y), we obtain,
$\frac{3}{2}h(X)+\frac{3}{2}h(Y)-\log C+I(X+Y;Y)\leq h(X)+2h(Y)-I(X;Y),$
and, rearranging, gives (31). As we only combined the assumed bound (30) with equalities, reversing the argument immediately shows that (31) implies (30). The equivalence of (30) with (32) follows by symmetry. Combining (31) and (32) yields (33), which implies (34).   $\Box$
The reason why large additive energy $a(X,Y)$ implies $h(X)$ and $h(Y)$ are close is, in part, because our definition of what it means for $a(X,Y)$ to be large is symmetric in $X$ and $Y$.
In the asymmetric case where $h(X)$ and $h(Y)$ are significantly different, with $h(X)<h(Y)$, say, then in view of (25) we can instead consider $a(X,Y)$ to be large if $a(X,Y)\geq 2h(X)+h(Y)-\log C$. Similarly, since $h(X+Y)\geq h(Y|X)$, we can consider $h(X+Y)$ to be small if $h(X+Y)\leq h(Y|X)+\log C$. The following result states that, if $a(X,Y)$ is large in this sense, then $h(X+Y)$ is appropriately small, with a partial converse when $X$ and $Y$ are weakly dependent. This strengthens and generalizes [5, Proposition 8] to the continuous case. The proof of Corollary 3.6 is essentially identical to that of Corollary 3.4, so we omit it.
Corollary 3.6.
Let $X$ and $Y$ be continuous random variables and $C$ be a constant. Then
$a(X,Y)\geq 2h(X)+h(Y)-\log C,$
if and only if
$h(X+Y)\leq h(Y)+\log C-2I(X;Y)\leq h(Y|X)+\log C.$
Finally, the following result can be viewed as a version of the contrapositive of Corollary 3.3. It states that, if the entropy $h(X+Y)$ is large, then the entropic additive energy $a(X,Y)$ is small, again with a partial converse when $X$ and $Y$ are only weakly dependent. Corollary 3.7, which follows immediately from Lemma 3.2, strengthens [5, Proposition 10] and generalizes it to the continuous case.
Corollary 3.7.
Let $X$ and $Y$ be continuous random variables and $C$ be a constant. Then,
$h(X+Y)\geq h(X)+h(Y)+\log C,$
if and only if,
$a(X,Y)\leq h(X)+h(Y)+\log C-2I(X;Y).$
4 The BSG theorem for differential entropy
Here we establish an entropic version of an important result in additive combinatorics, known as the Balog–Szemerédi–Gowers (BSG) theorem [1, 6]. The BSG theorem further quantifies the intuition described earlier, that sets $A$ and $B$ with large additive energy $E(A,B)$ must contain sizeable additive structure. Specifically, it states that if $E(A,B)$ is large, then $A$ and $B$ must contain high-density subsets $A^{\prime}$ and $B^{\prime}$ such that $\left\lvert A^{\prime}+B^{\prime}\right\rvert$ is small. It can be viewed as a partial converse to the observation (12) that a small sumset $\left\lvert A+B\right\rvert$ implies a large additive energy $E(A,B)$. Note that a full converse does not hold in general; there exist sets $A$ and $B$ with large $E(A,B)$ but small $\left\lvert A+B\right\rvert$, so the conclusion of the BSG theorem is best possible, up to quantitative constants.
Theorem 4.1 (Combinatorial BSG theorem).
Let $A$ and $B$ be finite subsets of the same additive group, and suppose $E(A,B)\geq c\left\lvert A\right\rvert^{3/2}\left\lvert B\right\rvert^{3/2}$ for some constant $c$. Then there are positive constants $^{\prime}, c′′c^{\prime\prime}$, and $C$ depending only on $c$ such that there exist $A^{\prime}\subset A$ and $B^{\prime}\subset B$ with $\left\lvert A^{\prime}\right\rvert\geq c^{\prime}\left\lvert A\right\rvert$ and $\left\lvert B^{\prime}\right\rvert\geq c^{\prime\prime}\left\lvert B\right\rvert$ satisfying
$\left\lvert A^{\prime}+B^{\prime}\right\rvert\leq C\left\lvert A\right\rvert^{1/2}\left\lvert B\right\rvert^{1/2}.$
An entropic BSG theorem was established by Tao in [25, Theorem 3.1], and its differential entropy analog was proved in [12, Theorem 3.14]. More recently, Gowers, Green, Manners, and Tao [8, Lemma A.2] proved a similar result with better constants for discrete entropy. However, the most faithful entropy analog of the BSG theorem is due to Goh [5], which is also the only one stated in terms of entropic additive energy.
Since the natural probabilistic analog of restricting to a subset is conditioning, all entropic BSG theorems establish the existence of particular conditionings that lead to appropriate bounds. Specifically, suppose the additive energy $A(X,Y)$ is large in the sense of (20), and consider conditionally independent versions $(X_{1},Y_{1})$ and $(X_{2},Y_{2})$. Then Goh’s [5, Theorem 6] says that the conditional entropies $H(X_{1}|X+Y)$ and $H(Y_{2}|X+Y)$ are close to $H(X)$ and $H(Y)$, respectively, while the conditional entropy of the sum, $H(X_{1}+Y_{2}|X+Y)$, is small, again in the sense of (20).
In this section, we state and prove the differential entropy analog of Goh’s BSG theorem. Although the proof in [5] relies on functional submodularity and hence does not generalize to the case of differential entropy, our proof of Theorem 4.3 below applies verbatim to the discrete case as well. We begin by establishing a useful upper bound on the additive energy. The discrete version of Lemma 4.2 first appeared in [8].
Lemma 4.2.
Suppose $X$ and $Y$ are continuous random variables, and let $(X_{1},Y_{1})$ and $(X_{2},Y_{2})$ be conditionally independent versions of $(X,Y)$ given $X+Y$. Then
$\max\{h(X_{1}-X_{2}),h(X_{1}-Y_{2})\}\leq 2h(X)+2h(Y)-a(X,Y).$ (35)
Proof. We first prove the inequality for the $h(X_{1}-Y_{2})$ term. By our assumptions, $X_{1}+Y_{1}=X_{2}+Y_{2}$, so,
$X_{1}-Y_{2}$ $=(X_{1}+Y_{1})-Y_{1}-Y_{2}$
$=(X_{2}+Y_{2})-Y_{1}-Y_{2}$
$=X_{2}-Y_{1}.$
And since for arbitrary random variables $A,B$ and $C$ we always have, $h(A|B,C)\leq h(A|B)$, we can bound
$h(X_{1},Y_{1},X_{1}-Y_{2})+h(X_{1}-Y_{2})$ $=h(X_{1},X_{1}-Y_{2})+h(Y_{1}|X_{1},X_{1}-Y_{2})+h(X_{1}-Y_{2})$
$\leq h(X_{1},X_{1}-Y_{2})+h(Y_{1}|X_{1}-Y_{2})+h(X_{1}-Y_{2})$
$=h(X_{1},X_{1}-Y_{2})+h(Y_{1},X_{1}-Y_{2})$
$=h(X_{1},Y_{2})+h(Y_{1},X_{2}-Y_{1})$
$\leq h(X)+h(Y)+h(Y_{1},X_{2})$
$\leq 2h(X)+2h(Y).$
Thus,
$h(X_{1}-Y_{2})$ $\leq 2h(X)+2h(Y)-h(X_{1},Y_{1},X_{1}-Y_{2}),$
so it suffices to show
$h(X_{1},Y_{1},X_{1}-Y_{2})=a(X,Y).$
But this immediately follows from noting that
$h(X_{1},Y_{1},X_{1}-Y_{2})=h(X_{1},Y_{1},Y_{2})=h(X_{1},X_{2},X+Y)=a(X,Y).$
Similarly, for the $h(X_{1}-X_{2})$ term, we have
$h(X_{1},Y_{1},X_{1}-X_{2})+h(X_{1}-X_{2})$ $=h(X_{1},X_{1}-X_{2})+h(Y_{1}|X_{1},X_{1}-X_{2})+h(X_{1}-X_{2})$
$\leq h(X_{1},X_{2})+h(Y_{1}|X_{1}-X_{2})+h(X_{1}-X_{2})$
$=h(X_{1},X_{2})+h(Y_{1},X_{1}-X_{2})$
$=h(X_{1},X_{2})+h(Y_{1},Y_{2}-Y_{1})$
$=h(X_{1},X_{2})+h(Y_{1},Y_{2})$
$\leq 2h(X)+2h(Y).$
Thus it again suffices to show $h(X_{1},Y_{1},X_{1}-X_{2})=a(X,Y)$, which follows from observing
$h(X_{1},Y_{1},X_{1}-X_{2})=h(X_{1},Y_{1},X_{2})=h(X_{1},X_{2},X+Y)=a(X,Y).$
Combining the two bounds yields (35).   $\Box$
As alluded to in [8], Lemma 4.2 allows one to prove a better version of the entropic BSG theorem; the original entropic BSG was proved in the discrete case by Tao in [25, Theorem 3.1] and in the continuous case by Kontoyiannis and Madiman in [12, Theorem 3.14].
Theorem 4.3 (Entropic BSG theorem).
Suppose the continuous random variables $X$ and $Y$ satisfy, for some constant $C$,
$a(X,Y)\geq\frac{3}{2}h(X)+\frac{3}{2}h(Y)-\log C.$
Then, taking $(X_{1},Y_{1})$ and $(X_{2},Y_{2})$ to be conditionally independent versions of $(X,Y)$ given $X+Y$, we have
$h(X_{1}|X+Y)$ $\geq h(X)-2\log C,$ (36)
$h(Y_{2}|X+Y)$ $\geq h(Y)-2\log C,$ (37)
and, moreover, $X_{1}$ and $Y_{2}$ are conditionally independent given $X+Y$, and satisfy
$h(X_{1}+Y_{2}|X+Y)\leq\frac{1}{2}h(X)+\frac{1}{2}h(Y)+\log C.$ (38)
Proof. With $(X,Y),(X_{1},Y_{1}),(X_{2},Y_{2})$ as in the theorem statement, and using Lemma 3.1, we have
$h(X_{1}|X+Y)$ $=h(X_{1},X+Y)-h(X+Y)$
$=h(X_{1},Y_{1})-h(X+Y)$
$=a(X,Y)-h(X,Y)$
$\geq\frac{3}{2}h(X)+\frac{3}{2}h(Y)-\log C-h(X,Y)$
$\geq\frac{1}{2}h(X)+\frac{1}{2}h(Y)-\log C.$
We similarly obtain
$h(Y_{2}|X+Y)\geq\frac{1}{2}h(X)+\frac{1}{2}h(Y)-\log C.$
Summing these two bounds yields
$h(X_{1}|X+Y)+h(Y_{2}|X+Y)\geq h(X)+h(Y)-2\log C,$
from which it follows that
$h(X_{1}|X+Y)$ $\geq h(X)+h(Y)-h(Y_{2}|X+Y)-2\log C$
$\geq h(X)-2\log C.$
This proves (36), and (37) follows by symmetry.
Since $(X_{1}+Y_{2})-(X+Y)=X_{1}-X_{2}$, we have
$h(X_{1}+Y_{2}|X+Y)$ $=h(X_{1}+Y_{2},X+Y)-h(X+Y)$
$=h(X_{1}-X_{2},X+Y)-h(X+Y)$
$=h(X_{1}-X_{2}|X+Y)$
$\leq h(X_{1}-X_{2}),$
and using Lemma 4.2,
$h(X_{1}+Y_{2}|X+Y)$ $\leq 2h(X)+2h(Y)-a(X,Y)$
$\leq\frac{1}{2}h(X)+\frac{1}{2}h(Y)+\log C.$
This gives (38) and completes the proof.   $\Box$
As observed by Goh [5], the assumptions of the BSG theorems in [25, Theorem 3.1] and [12, Theorem 3.14] combined with Corollary 3.3 imply that
$a(X,Y)\geq\frac{3}{2}h(X)+\frac{3}{2}h(Y)-3\log K,$
and Theorem 4.3 then gives (36)–(38) with error terms $6\log K$, $6\log K$, and $3\log K$, respectively. Meanwhile, the analogous statements in the BSG statements of [25] and [12] have error terms $\log K$, $\log K$, and $7\log K$, respectively. Thus, Theorem 4.3 has worse bounds on the conditional entropies of $X_{1}$ and $Y_{2}$, but better bounds on the “main” inequality (38) on the conditional entropy of the conditionally independent sum $X_{1}+Y_{2}$.
5 Stability of large discrete entropic doubling
Before continuing with our development of differential entropy inequalities, in this section we temporarily return to the case of discrete entropy. For a discrete random variable $X$ taking values in a subset $A$ of an additive group $G$, we define the doubling constant $s(X)$ in analogy with the combinatorial doubling constant $s(A)$ in (8), via,
$s(X)=H(X+X^{\prime})-H(X),$ (39)
where $X,X^{\prime}$ are i.i.d. Clearly, $s(X)\geq 0$, and Tao [25] showed that $s(X)$ is small if and only if $X$ is approximately uniformly distributed on a generalized arithmetic progression. Here, we examine the behaviour of the doubling constant $s(X)$ at its other extreme. We note that $s(X)$ is always bounded above by $H(X)$, and we show that $s(X)$ is close to $H(X)$ if and only if $X$ is approximately supported on a Sidon set. Recall that $A$ is a Sidon set if, whenever $a+b=c+d$ for $a,b,c,d\in A$, we necessarily have $\left\{a,b\right\}=\left\{c,d\right\}$.
Let $X$ be a discrete random variable with probability mass function $P$ on $A$. Goh [5] observed that, if $X$ is supported on a Sidon set, i.e., if $\{a\in G:P(a)>0\}$ is a Sidon set, then $s(X)\geq H(X)-\log 2$. Our first result identifies a tighter upper bound to $s(X)$ than the trivial bound $H(X)$, and shows that it is achieved if and only if $X$ is supported on a Sidon set.
Lemma 5.1.
If $X$ is discrete random variable with values in a subset $A$ of an additive group $G$, and with probability mass function $P$, then,
$s(X)\leq H(X)-(\log 2)\left(1-\sum_{a\in A}P(a)^{2}\right),$ (40)
with equality if and only if $X$ is supported on a Sidon set.
Proof. By the definition of $s(X)$,
$H(X)-s(X)$ $=2H(X)-H(X+X^{\prime})$
$=H(X,X^{\prime}|X+X^{\prime})$
$=\sum_{a,b\in A}P(a)P(b)\log\Big(\frac{\sum_{c\in A}P(c)P(a+b-c)}{P(a)P(b)}\Big).$
This expression together with the observation that
$\sum_{c\in A}P(c)P(a+b-c)\geq P(a)P(b)+P(b)P(a)=2P(a)P(b),$ (41)
whenever $a\neq b$, immediately imply (40).
Conversely, in order to have equality in (40) we must have equality in (41), which can only happen if any possible value $a+b$ of the sum $X+X^{\prime}$ that has nonzero probability can only be achieved by $(X,X^{\prime})=(a,b)$ or $(X^{\prime},X)=(a,b)$; this implies that $X$ is supported on a Sidon set.   $\Box$
We now prove a stability result, showing that if $s(X)$ is close to its upper bound in (40), then $X$ is close to being supported on a Sidon set.
Proposition 5.2.
Let $X$ be a discrete random variable supported on a finite subset $A$ of an additive group $G$, and with probability mass function $P$. If
$s(X)\geq H(X)-(\log 2)\left(1-\sum_{a\in A}P(a)^{2}\right)-C,$ (42)
for some constant $C\geq 0$, then there exists a Sidon set $B\subset A$ such that
${\mathbb{P}}(X\in B)\geq 1-\frac{C}{p_{*}(\log 2)},$ (43)
where $p_{*}=\min\{P(a):a\in A\}>0$.
Proof. For $a,b\in A$, let
$R(a,b)=\log\Big(\frac{\sum_{c\in A}P(c)P(a+b-c)}{P(a)P(b)}\Big)-(\log 2){\mathbb{I}}_{\{a\neq b\}}\geq 0.$
To see that $R(a,b)$ is nonnegative note that, for any $a,b\in A$,
$\sum_{c\in A}P(c)P(a+b-c)\geq\frac{2P(a)P(b)}{1+{\mathbb{I}}_{\{a=b\}}}.$
Then our assumption is equivalent to the inequality ${\mathbb{E}}[R(X,X^{\prime})]\leq C$, where $X$ and $X^{\prime}$ are i.i.d., and by Markov’s inequality, we have
${\mathbb{P}}(R(X,X^{\prime})\geq\log 2)\leq\frac{C}{\log 2}.$ (44)
We construct a subset $D\subset A$ as follows. Let $I$ be the set of unordered pairs $\{a,b\}$ with $R(a,b)\geq\log 2$, noting that $R(a,b)=R(b,a)$ so this is a well-defined set. For any $\{a,b\}\in I$, if $P(a)\neq P(b)$, then include in $D$ the element with the smaller probability; otherwise, if $P(a)=P(b)$, include in $D$ one of the two, choosing arbitrarily. Let $B=A\setminus D$.
We first prove (43). By (44) and the definition of $D$, we have
${\mathbb{P}}(X\in D)$ $\leq\sum_{\{a,b\}\in I}\min\{P(a),P(b)\}$
$\leq\sum_{\{a,b\}\in I}\frac{P(a)P(b)}{p_{*}}\left(1+{\mathbb{I}}_{\{a\neq b\}}\right)$
$=\frac{1}{p_{*}}{\mathbb{P}}(R(X,X^{\prime})\geq\log 2)$
$\leq\frac{C}{p_{*}(\log 2)},$
which is equivalent to (43).
It remains to show $B$ is a Sidon set. Notice that for all $a,b\in B$, we have $R(a,b)<\log 2$. Suppose $B$ is not a Sidon set, i.e., that there are $a_{1},b_{1},a_{2},b_{2}\in B$ are such that $a_{1}+b_{1}=a_{2}+b_{2}$ and $\{a_{1},b_{1}\}\neq\{a_{2},b_{2}\}$. This implies that,
$\sum_{c\in A}P(c)P(a_{1}+b_{1}-c)\geq\frac{2P(a_{1})P(b_{1})}{1+{\mathbb{I}}_{\{a_{1}=b_{1}\}}}+\frac{2P(a_{2})P(b_{2})}{1+{\mathbb{I}}_{\{a_{2}=b_{2}\}}}.$ (45)
Now consider two cases. If
$\frac{P(a_{1})P(b_{1})}{1+{\mathbb{I}}_{\{a_{1}=b_{1}\}}}\leq\frac{P(a_{2})P(b_{2})}{1+{\mathbb{I}}_{\{a_{2}=b_{2}\}}},$
then (45) and the definition of $R(a,b)$ imply that $R(a_{1},b_{1})\geq\log 2$, while if the reverse inequality holds, we have $R(a_{2},b_{2})\geq\log 2$, a contradiction either way. Hence, $B$ is a Sidon set, and the proof is complete.   $\Box$
The following two examples demonstrate that the assumption that $X$ takes on only finitely many values – equivalently, that $p_{*}$ is strictly positive – cannot be removed. In fact, the dependence of the bound (43) on $p_{*}$ cannot be entirely avoided.
Example 5.3.
Consider an $N$-element set $A\subset G$ that has exactly one violation to the Sidon set condition, that is, exactly one nontrivial solution to $a_{1}+b_{1}=a_{2}+b_{2}$, and suppose also that, in that case, both $a_{1}\neq b_{1}$ and $a_{2}\neq b_{2}$. Such an $A$ can be easily constructed in $G={\mathbb{Z}}$, for example. Let $X$ be the uniform distribution on $A$. Then we see that
$s(X)\geq H(X)-(\log 2)\left(1-\sum_{a\in A}P(a)^{2}\right)-\frac{4\log 2}{N^{2}},$
so the assumption of Proposition 5.2 holds with $C=\frac{4\log 2}{N^{2}}$, while for any Sidon set $B\subset G$ we have
${\mathbb{P}}(X\in B)\leq\frac{N-1}{N}=1-\frac{1}{N}=1-\frac{C}{4p_{*}(\log 2)}.$
Example 5.4.
A more striking example is as follows. Consider the $4N$-element subset $A\subset{\mathbb{Z}}$ given by
$A=\bigcup_{k=0}^{N-1}\big\{10^{k},2\cdot 10^{k},4\cdot 10^{k},5\cdot 10^{k}\big\}.$
It is easy to see that the only violations to the Sidon set condition are of the form
$10^{k}+5\cdot 10^{k}=2\cdot 10^{k}+4\cdot 10^{k}.$
Thus, letting $X$ be the uniform distribution on $A$, we have
$s(X)\geq H(X)-(\log 2)\left(1-\sum_{a\in A}P(a)^{2}\right)-\frac{\log 2}{4N},$
yet any Sidon set $B\subset{\mathbb{Z}}$ can contain at most 3 of the 4 elements of $\{10^{k},2\cdot 10^{k},4\cdot 10^{k},5\cdot 10^{k}\}$, which implies
${\mathbb{P}}(X\in B)\leq\frac{3}{4}=1-\frac{1}{4}.$
In other words, as $N\to\infty$, the condition of Proposition 5.2 holds with $C=\frac{\log 2}{4N}\to 0$, while for any Sidon set $B$ the probability ${\mathbb{P}}(X\in B)$ stays bounded away from 1. Still, this does not violate the result of Proposition 5.2, because here we have $\frac{C}{p_{*}(\log 2)}=1$.
Proposition 5.2 and the two examples above raise the following question.
Open problem. It would be interesting to determine whether there exists a function $f(C,D)\geq 0$ for $C,D\geq 0$, with $f(C,D)\to 0$ as $C\to 0$ for any fixed $D\geq 0$, such that the following holds: If $X$ is a discrete random variable with values in $A$ and with $H(X)\leq D$ and satisfying condition (42) of Proposition 5.2 with constant $C$, then there exists a Sidon set $B\subset A$ such that ${\mathbb{P}}(X\in B)\geq 1-f(C,D).$
Finally, we return to the obvious upper bound $s(X)\leq H(X)$. Lemma 5.1 shows that this is achieved if and only if $X$ is deterministic. Our last result in this section is a stability statement, showing that if $s(X)$ is close to $H(X)$, then $X$ is close to being deterministic.
Proposition 5.5.
Let $X$ be a discrete random variable with values in a subset $A$ of an additive group $G$, and with probability mass function $P$. If
$s(X)\geq H(X)-\varepsilon$
for some constant $\varepsilon>0$, then
$p^{*}:=\max_{a\in A}P(a)\geq 1-\frac{\varepsilon}{\log 2}.$
Proof. As
$\sum_{a\in A}P(a)^{2}\leq p^{*}\sum_{a\in A}P(a)=p^{*},$
by Lemma 5.1 we have
$1-p^{*}\leq 1-\sum_{a\in A}P(a)^{2}\leq\frac{H(X)-s(X)}{\log 2}\leq\frac{\varepsilon}{\log 2},$
which gives the claimed inequality.   $\Box$
6 Differential entropy bounds for products
In this section we describe how all the additive entropy inequalities for sums of continuous random variables discussed in Section 2.3 have natural and easy to derive multiplicative analogs.
6.1 Multiplicative entropy
Let $X$ be an arbitrary continuous random variable. Since ${\mathbb{P}}(X=0)=0$, we can think of $X$ as taking values in the group ${\mathbb{R}}^{\times}:={\mathbb{R}}\setminus\{0\}$ equipped with the multiplication operation. Recall that $({\mathbb{R}}^{\times},\cdot)$ is an abelian group with Haar measure $\rho$ given by its density $\frac{d\rho}{d\lambda}(x)=\frac{1}{|x|}$, $x\in{\mathbb{R}}^{\times}$, with respect to Lebesgue measure $\lambda$. If $X$ has law $\mu$ with density $f=d\mu/d\lambda$ with respect to Lebesgue measure, then it also has density
$g(x)=\frac{d\mu}{d\rho}(x)=|x|f(x),\quad x\in{\mathbb{R}}^{\times},$
with respect to Haar measure.
In [16] it was observed that a number of entropy inequalities for sums, including those mentioned in Section 2.3, also hold for random variables taking values in arbitrary locally compact, Polish, abelian groups (G,)(G,*), with addition replaced by the group operation *, and with differential entropy replaced by its obvious analog,
$\widetilde{h}(X)=-\int g\log g\,d\rho,$
where $\rho$ denotes the Haar measure on the Borel $\sigma$-algebra of $G$. In the case of $({\mathbb{R}}^{\times},\cdot)$, we call $\widetilde{h}(X)$ the multiplicative entropy of a continuous random variable $X$, and observe that, if $X$ has density $f$ with respect to Lebesgue measure, then $\widetilde{h}(X)$ and the usual differential entropy $h(X)$ are related via
$\widetilde{h}(X)=-\int g\log g\,d\rho=-\int\left\lvert x\right\rvert f(x)\log(\left\lvert x\right\rvert f(x))\,\frac{dx}{\left\lvert x\right\rvert}=h(X)-{\mathbb{E}}[\log\left\lvert X\right\rvert$]. (46)
The joint multiplicative entropy of a finite collection of continuous random variables is defined in the obvious way, and we similarly define the conditional multiplicative entropy, the multiplicative mutual information, and so on.
Here and in Section 6.2 we note a number of “multiplicative” entropy inequalities that we will find useful in Section 7. All these results follow from the correspondence noted in [16] and, except for the simple computations in Lemmas 6.1 and 6.2, they are stated without proof. In all of our results, we implicitly assume that any differential entropies and multiplicative entropies appearing in the statement exist and are finite. Nevertheless, for the sake of clarity, we usually explicitly state the integrability assumptions required for the multiplicative entropy to be finite.
Just as differential entropy is translation invariant, $h(X+c)=h(X)$, multiplicative entropy is scale invariant:
Lemma 6.1.
For any continuous random variable $X$ such that $\log|X|$ is integrable,
$h(1/X)=h(X)-2{\mathbb{E}}[\log\left\lvert X\right\rvert$],
or, equivalently,
$\widetilde{h}(1/X)=\widetilde{h}(X).$
Proof. Let $f_{X},f_{1/X}$ denote the densities of $X$ and $1/X$ with respect to Lebesgue measure. Since $f_{1/X}(y)=(1/y^{2})f_{X}(y)$, $y\neq 0$, we have
$h(1/X)$ $=-\int_{\mathbb{R}}f_{1/X}(y)\log f_{1/X}(y)\,dy$
$=-\int_{\mathbb{R}}\frac{1}{y^{2}}f_{X}(1/y)\log\left(\frac{1}{y^{2}}f_{X}(1/y)\right)\,dy$
$=-\int_{\mathbb{R}}f_{X}(x)\log(x^{2}f_{X}(x))\,dx$
$=h(X)-2{\mathbb{E}}[\log\left\lvert X\right\rvert$],
where we note that the reversal of the range of integration “cancels out” with the negative sign from the change of variables. The second identity follows from (46).   $\Box$
Lemma 6.2.
Let $(X_{1},\ldots,X_{n},Y)$ be a continuous random vector, such that $\log|Y|$ is integrable. Then, for all integers $k_{1},k_{2},\dots,k_{n}$,
$h(X_{1}Y^{k_{1}},X_{2}Y^{k_{2}},\dots,X_{n}Y^{k_{n}}|Y)=h(X|Y)+K{\mathbb{E}}[\log\left\lvert Y\right\rvert$], (47)
where $X=(X_{1},\ldots,X_{n})$ and $K=k_{1}+\cdots+k_{n}$. Equivalently,
$\widetilde{h}(X_{1}Y^{k_{1}},X_{2}Y^{k_{2}},\dots,X_{n}Y^{k_{n}}|Y)=\widetilde{h}(X|Y).$
Proof. Let $Z=(X_{1}Y^{k_{1}},X_{2}Y^{k_{2}},\dots,X_{n}Y^{k_{n}})$. Using the obvious notation for joint and conditional densities and writing $z^{\prime}=(z_{1}/y^{k_{1}},z_{2}/y^{k_{2}},\dots,z_{n}/y^{k_{n}})$, we have
$h(Z,Y)$ $=-\int_{\mathbb{R}}\int_{{\mathbb{R}}^{n}}f_{Y}(y)f_{Z|Y}(z|y)\log(f_{Y}(y)f_{Z}(z|y))\,dzdy$
$=-\int_{\mathbb{R}}\int_{{\mathbb{R}}^{n}}\frac{1}{\left\lvert y\right\rvert^{K}}f_{Y}(y)f_{X|Y}\left(z^{\prime}\middle|y\right)\log\left(\frac{1}{\left\lvert y\right\rvert^{K}}f_{Y}(y)f_{X|Y}\left(z^{\prime}\middle|y\right)\right)\,|y|^{K}dz^{\prime}dy$
$=-\int_{\mathbb{R}}\int_{{\mathbb{R}}^{n}}f_{Y}(y)f_{X|Y}(x|y)\log\left(\frac{1}{\left\lvert y\right\rvert^{K}}f_{Y}(y)f_{X|Y}(x|y)\right)\,dxdy$
$=h(Y)+h(X|Y)+K\cdot{\mathbb{E}}[\log\left\lvert Y\right\rvert$],
which rearranges to the desired identity (47). The second identity follows from (46).   $\Box$
The following is the multiplicative version of Ruzsa’s triangle inequality.
Proposition 6.3 (Multiplicative Ruzsa triangle inequality).
Let $X$, $Y$, and $Z$ be independent continuous random variables such that $\log\left\lvert X\right\rvert$, $\log\left\lvert Y\right\rvert$, and $\log\left\lvert Z\right\rvert$ are integrable. Then
$h(X/Z)\leq h(X/Y)+h(Y/Z)-h(Y)+{\mathbb{E}}[\log\left\lvert Y\right\rvert$],
or, equivalently,
$\widetilde{h}(X/Z)\leq\widetilde{h}(X/Y)+\widetilde{h}(Y/Z)-\widetilde{h}(Y).$
Similarly, the multiplicative analog of the submodularity-for-sums bound in (22) is:
Proposition 6.4.
Let $X$, $Y$, and $Z$ be independent, continuous random variables, such that $\log\left\lvert X\right\rvert$, $\log|Y|$ and $\log|Z|$ are integrable. Then
$h(XYZ)+h(Y)\leq h(XY)+h(YZ),$
or, equivalently,
$\widetilde{h}(XYZ)+\widetilde{h}(Y)\leq\widetilde{h}(XY)+\widetilde{h}(YZ).$
Next, we define the multiplicative Ruzsa distance between two continuous random variables $X$ and $Y$, where both $\log\left\lvert X\right\rvert$ and $\log\left\lvert Y\right\rvert$ are integrable, in analog with the additive version in (21), as,
$\widetilde{d}(X,Y)$ $=h(X^{\prime}/Y^{\prime})-\frac{1}{2}h(X)-\frac{1}{2}h(1/Y)-\frac{1}{2}{\mathbb{E}}[\log\left\lvert X\right\rvert]+\frac{1}{2}{\mathbb{E}}[\log\left\lvert Y\right\rvert$]
$=\widetilde{h}(X^{\prime}/Y^{\prime})-\frac{1}{2}\widetilde{h}(X)-\frac{1}{2}\widetilde{h}(Y),$
where $X^{\prime}\sim X$ and $Y^{\prime}\sim Y$ are independent.
Note that the regular Ruzsa distance is translation invariant, $d(X+c_{1},Y+c_{2})=d(X,Y)$, while the multiplicative Ruzsa distance is scale invariant, $\widetilde{d}(c_{1}X,c_{2}Y)=\widetilde{d}(X,Y)$.
6.2 The multiplicative doubling constant
The doubling constant of a discrete random variable $X$ was defined in (39) as $s(X)=H(X+X^{\prime})-H(X)$, with $X,X^{\prime}$ being i.i.d. Similarly, we define the doubling constant $\sigma(X)$ and the difference constant $\delta(X)$ of a continuous random variable $X$ as,
$\sigma(X)=h(X+X^{\prime})-h(X),\qquad\delta(X)=h(X-X^{\prime})-h(X),$
where $X,X^{\prime}$ are i.i.d. In this notation, the doubling-difference inequality for differential entropy described in Section 2.3, states that
$\frac{1}{2}\leq\frac{\sigma(X)}{\delta(X)}\leq 2.$ (48)
The natural multiplicative analog of $\sigma(X)$ and $\delta(X)$ are,
$\widetilde{\sigma}(X):=h(XX^{\prime})-h(X)-{\mathbb{E}}[\log\left\lvert X\right\rvert]=\widetilde{d}(X,1/X),$
and,
$\widetilde{\delta}(X)=h(X/X^{\prime})-h(X)+{\mathbb{E}}[\log\left\lvert X\right\rvert]=\widetilde{d}(X,X),$
where $X,X^{\prime}$ are i.i.d. The following “square-quotient inequality” is the multiplicative version of the doubling-difference inequality.
Theorem 6.5 (Square-quotient inequality).
Let $X$ be a continuous random variable such that $\log\left\lvert X\right\rvert$ is integrable. Then
$\frac{1}{2}\leq\frac{\widetilde{\sigma}(X)}{\widetilde{\delta}(X)}\leq 2,$
or, equivalently,
$\frac{1}{2}\widetilde{d}(X,X)\leq\widetilde{d}(X,1/X)\leq 2\widetilde{d}(X,X).$
Finally, we observe that the same inductive application of Proposition 6.4 as was done for the additive case in [16], immediately gives the following multiplicative version of the Plünnecke–Ruzsa inequality.
Theorem 6.6 (Multiplicative Plünnecke–Ruzsa inequality).
Let $X$ and $Y_{1},Y_{2},\dots,Y_{n}$ be independent, continuous random variables, such that $\log\left\lvert X\right\rvert$ and all $\log|Y_{i}|$, $1\leq i\leq n$, are integrable. Suppose there are finite constants $K_{1},K_{2},\dots,K_{n}$ satisfying $h(XY_{i})\leq h(X)+\log K_{i}$ for each $i$. Then:
$h(XY_{1}Y_{2}\cdots Y_{n})\leq h(X)+\log(K_{1}K_{2}\cdots K_{n}).$
7 Differential entropy bounds for sum-product combinations
The main goal of this section is to establish two general versions of the Plünnecke–Ruzsa inequality that involve both sums and products of i.i.d. random variables. The first one, in Theorem 7.6, applies to real-valued continuous random variables and the second one, in Theorem 7.7, holds for discrete random variables taking values in an arbitrary integral domain.
We begin by proving the following inequality for sums and products of an arbitrary triple $(X,Y,Z)$ of continuous random variables. It is the continuous version of a corresponding inequality proved for discrete entropy by Máthé and O’Regan in [18, Proposition 4.1]. Proposition 7.1 will be used in the proof of our first version of the ring Plünnecke–Ruzsa inequality in Theorem 7.2.
Proposition 7.1.
Let $X,Y,Z$ be continuous random variables such that $\log\left\lvert X\right\rvert$ is integrable. Then
$h(X(Y+Z))+h(X,Y,Z)\leq h(X,Y+Z)+h(XY,XZ)-{\mathbb{E}}[\log\left\lvert X\right\rvert$]. (49)
In particular, if $X$, $Y$, and $Z$ are i.i.d., then
$h(X(Y\pm Z))\leq 2h(XY)+h(X\pm Y)-2h(X)-{\mathbb{E}}[\log\left\lvert X\right\rvert$], (50)
where the inequality holds with either choice of signs, $(+,+)$ or (,)(-,-).
Proof. Recall that $h(aW)=h(W)+\log|a|$ for any continuous random variable $W$ and any nonzero constant $a$. Then we note that we may express
$I(X(Y+Z);X)$ $=h(X(Y+Z))-h(X(Y+Z)|X)$
$=h(X(Y+Z))-h(Y+Z|X)-{\mathbb{E}}[\log\left\lvert X\right\rvert$]
$=h(X(Y+Z))-h(X,Y+Z)+h(X)-{\mathbb{E}}[\log\left\lvert X\right\rvert$],
and
$I(XY,XZ;X)$ $=h(XY,XZ)-h(XY,XZ|X)$
$=h(XY,XZ)-h(Y,Z|X)-2{\mathbb{E}}[\log\left\lvert X\right\rvert$]
$=h(XY,XZ)-h(X,Y,Z)+h(X)-2{\mathbb{E}}[\log\left\lvert X\right\rvert$].
By the data processing inequality, we have
$I(X(Y+Z);X)\leq I(XY,XZ;X),$
which rearranges to (49).
Assuming $X$, $Y$, and $Z$ are i.i.d., applying (49) to $X$, $Y$, and $\pm Z$, and using the simple fact that
$h(XY,XZ)\leq h(XY)+h(XZ)=2h(XY).$
yields (50).   $\Box$
7.1 Ring Plünnecke–Ruzsa inequality
We begin by first establishing the following simple version of a “ring” Plünnecke–Ruzsa inequality, which provides an upper bound to h(XYZW)h(XY-ZW) when $X,Y,Z,W$ are i.i.d. Recall the definitions of the doubling constant $\sigma(X)$ and of the associated constants $\delta(X)$ and $\widetilde{\sigma}(X)$ from Section 6.2.
Theorem 7.2 (Ring Plünnecke–Ruzsa).
If $X,Y,Z,W$ are i.i.d. continuous random variables such that $\log\left\lvert X\right\rvert$ is integrable, then,
$h(XY-ZW)$ $\leq 5h(XY)+2h(X+Y)-6h(X)-4{\mathbb{E}}[\log\left\lvert X\right\rvert$] (51)
$=h(XY)+4\widetilde{\sigma}(X)+2\sigma(X),$
and
$h(XY+ZW)$ $\leq 5h(XY)+h(X+Y)+h(X-Y)-6h(X)-4{\mathbb{E}}[\log\left\lvert X\right\rvert$] (52)
$=h(XY)+4\widetilde{\sigma}(X)+\sigma(X)+\delta(X).$
In addition,
$h(XY+ZW)\leq 5h(XY)+3h(X\pm Y)-7h(X)-4{\mathbb{E}}[\log\left\lvert X\right\rvert$]. (53)
For the proof we will need the following bound.
Lemma 7.3.
For any three continuous random variables $X,Y,Z$,
$h(X\pm Y)+h(X,Y,Z)\leq h(X,Y)+h(X\mp Z,Y+Z),$ (54)
where the inequality holds with either choice of signs, (+,)(+,-) or (,+)(-,+).
Proof. Note that
$I(X\pm Y;Y+Z)$ $=h(X\pm Y)+h(Y+Z)-h(X\pm Y,Y+Z)$
$=h(X\pm Y)+h(Y+Z)-h(X\mp Z,Y+Z)$
and
$I(X,Y;Y+Z)$ $=h(X,Y)+h(Y+Z)-h(X,Y,Y+Z)$
$=h(X,Y)+h(Y+Z)-h(X,Y,Z).$
By the data processing inequality, we have
$I(X\pm Y;Y+Z)\leq I(X,Y;Y+Z),$
which rearranges to (54).   $\Box$
Proof of Theorem 7.2. Let $V=(XY,ZW,XW)$. Note that by Lemma 6.2,
$h(V|X)$ $=h(Y,ZW,W|X)+2{\mathbb{E}}[\log\left\lvert X\right\rvert$]
$=h(Y)+h(ZW,W)+2{\mathbb{E}}[\log\left\lvert X\right\rvert$]
$=2h(X)+h(ZW|W)+2{\mathbb{E}}[\log\left\lvert X\right\rvert$]
$=2h(X)+h(Z|W)+3{\mathbb{E}}[\log\left\lvert X\right\rvert$]
$=3h(X)+3{\mathbb{E}}[\log\left\lvert X\right\rvert$].
Also,
$h(X|V)$ $\leq h(X|XY)$
$=h(X,XY)-h(XY)$
$=h(X)+h(Y|X)-h(XY)+{\mathbb{E}}[\log\left\lvert X\right\rvert$]
$=2h(X)-h(XY)+{\mathbb{E}}[\log\left\lvert X\right\rvert$],
which implies the lower bound
$h(V)$ $=h(V|X)+h(X)-h(X|V)$
$\geq 2h(X)+h(XY)+2{\mathbb{E}}[\log\left\lvert X\right\rvert$]. (55)
Applying Lemma 7.3 to the three random variables $XY,ZW,XW$, we obtain
$h(XY\pm ZW)+h(V)$ $\leq h(XY,ZW)+h(XY\mp XW,ZW+XW).$
This, combined with the lower bound in (55), yields
$h(XY\pm ZW)+2h(X)+h(XY)+2{\mathbb{E}}[\log\left\lvert X\right\rvert$] $\leq h(XY,ZW)+h(XY\mp XW,ZW+XW)$
$\leq h(XY,ZW)+h(XY\mp XW)+h(ZW+XW),$
and using the fact that $X,Y,Z,W$ are i.i.d. and rearranging,
$h(XY\pm ZW)\leq h(XY)+h(X(Y+Z))+h(X(Y\mp Z))-2h(X)-2{\mathbb{E}}[\log\left\lvert X\right\rvert$].
Finally, applying Proposition 7.1 yields (51) and (52), and the doubling-difference inequality (48) combined with (52), yields (53).   $\Box$
A more general, inductive version of the argument in the last proof, allows us to prove the following technical result, which is the key step in establishing the general ring Plünnecke–Ruzsa inequality below.
Proposition 7.4.
Suppose the continuous random variables $X,X_{1},X_{2},\dots,X_{n},Y_{1},Y_{2},\dots,Y_{n}$ are i.i.d., and that $\log\left\lvert X\right\rvert$ is integrable. Then
$h(X_{1}X_{2}\cdots X_{n}+Y_{1}Y_{2}\cdots Y_{n})$ $\leq 3h(X_{1}X_{2}\cdots X_{n})+2\sum_{k=2}^{n}h(X_{1}X_{2}\cdots X_{k})+(n-1)h(X_{1}-Y_{1})$
$\quad+h(X_{1}+Y_{1})-3nh(X)-(n+2)(n-1){\mathbb{E}}[\log\left\lvert X\right\rvert$].
Proof. We prove the result by induction. The case $n=1$ is trivial, and the case $n=2$ follows from Theorem 7.2. For the inductive step, assume $n\geq 3$. Let $A=X_{1}X_{2}\cdots X_{n-1}$ and let $B=Y_{1}Y_{2}\cdots Y_{n-1}$. We mimic the proof of Theorem 7.2, except using the less symmetric random vector $Z=(AX_{n},BY_{n},AY_{n})$ in place of $V$. By Lemma 6.2,
$h(Z|A)$ $=h(X_{n},BY_{n},Y_{n})+2(n-1){\mathbb{E}}[\log\left\lvert X\right\rvert$]
$=2h(X)+h(BY_{n}|Y_{n})+2(n-1){\mathbb{E}}[\log\left\lvert X\right\rvert$]
$=2h(X)+h(A)+(2n-1){\mathbb{E}}[\log\left\lvert X\right\rvert$].
We also have,
$h(A|Z)$ $\leq h(A|AX_{n})$
$=h(A)+h(AX_{n}|A)-h(AX_{n})$
$=h(A)+h(X_{n})-h(AX_{n})+(n-1){\mathbb{E}}[\log\left\lvert X\right\rvert$].
Thus,
$h(Z)$ $=h(Z|A)+h(A)-h(A|Z)$
$\geq h(AX_{n})+h(A)+h(X)+n{\mathbb{E}}[\log\left\lvert X\right\rvert$]. (56)
Meanwhile, using Lemma 7.3 applied to $Z$, and Proposition 7.1, we obtain,
$h(AX_{n}+BY_{n})+h(Z)$ $\leq h(AX_{n},BY_{n})+h(A(X_{n}-Y_{n}),(A+B)Y_{n})$
$\leq 2h(AX_{n})+h(A(X_{n}-Y_{n}))+h((A+B)Y_{n})$
$\leq 2h(AX_{n})+h(A,X_{n}-Y_{n})-h(A,X_{n},Y_{n})+h(AX_{n},AY_{n})$
$\quad-(n-1){\mathbb{E}}[\log\left\lvert X\right\rvert]+h(A+B,Y_{n})-h(A,B,Y_{n})$
$\quad+h(AY_{n},BY_{n})-{\mathbb{E}}[\log\left\lvert X\right\rvert$]
$\leq 6h(AX_{n})+h(A+B)+h(X_{n}-Y_{n})-2h(X)-2h(A)-n{\mathbb{E}}[\log\left\lvert X\right\rvert$],
and using the lower bound (56) on $h(Z)$, gives,
$h(AX_{n}+BY_{n})$ $\leq 5h(AX_{n})+h(A+B)+h(X_{n}-Y_{n})-3h(X)-3h(A)-2n{\mathbb{E}}[\log\left\lvert X\right\rvert$].
Finally, applying the inductive hypothesis to $h(A+B)$,
$h(AX_{n}+BY_{n})$ $\leq 5h(AX_{n})+3h(A)+2\sum_{k=2}^{n-1}h(X_{1}X_{2}\cdots X_{k})+(n-2)h(X_{1}-Y_{1})+h(X_{1}+Y_{1})$
$\quad-3(n-1)h(X)-(n+1)(n-2){\mathbb{E}}[\log\left\lvert X\right\rvert]+h(X_{n}-Y_{n})-3h(X)-3h(A)$
$\quad-2n{\mathbb{E}}[\log\left\lvert X\right\rvert$]
$=3h(AX_{n})+2\sum_{k=2}^{n}h(X_{1}X_{2}\cdots X_{k})+(n-1)h(X_{1}-Y_{1})+h(X_{1}+Y_{1})-3nh(X)$
$\quad-(n+2)(n-1){\mathbb{E}}[\log\left\lvert X\right\rvert$],
which is exactly the claimed result.   $\Box$
Applying the multiplicative Plünnecke–Ruzsa inequality, Theorem 6.6, to Proposition 7.4, immediately yields the following corollary.
Corollary 7.5.
Suppose the continuous random variables $X,X_{1},X_{2},\dots,X_{n},Y_{1},Y_{2},\dots,Y_{n}$ are i.i.d., and that $\log\left\lvert X\right\rvert$ is integrable. Then
$h(X_{1}X_{2}\cdots X_{n}+Y_{1}Y_{2}\cdots Y_{n})$ $\leq h(X_{1}X_{2}\cdots X_{n})+(n+2)(n-1)\big[h(X_{1}Y_{1})-h(X)-{\mathbb{E}}[\log\left\lvert X\right\rvert$$\big]$
$\quad+(n-1)\big[h(X_{1}-Y_{1})-h(X)\big]+h(X_{1}+Y_{1})-h(X)$ (57)
$=h(X_{1}X_{2}\cdots X_{n})+(n+2)(n-1)\widetilde{\sigma}(X)+(n-1)\delta(X)+\sigma(X).\quad\quad$
Finally, if we apply the additive Plünnecke–Ruzsa inequality (23) of [12] to (57), and then use the multiplicative Plünnecke–Ruzsa inequality, Theorem 6.6, we obtain the following general ring Plünnecke–Ruzsa inequality. This is the main result of this section.
Theorem 7.6 (General ring Plünnecke–Ruzsa inequality).
Suppose that continuous random variables $\{X_{ij}\;;\;1\leq i\leq m,1\leq j\leq n\}$ are i.i.d. and distributed as $X$. Assuming $\log\left\lvert X\right\rvert$ is integrable, we have
$h\left(\sum_{i=1}^{m}\prod_{j=1}^{n}X_{i,j}\right)$ $\leq h(X_{1,1}X_{1,2}\cdots X_{1,n})+(m-1)\big[(n+2)(n-1)\widetilde{\sigma}(X)+(n-1)\delta(X)+\sigma(X)\big]$
$\leq h(X)+\big((m-1)(n+2)+1\big)(n-1)\widetilde{\sigma}(X)+(m-1)(n-1)\delta(X)+(m-1)\sigma(X)$
$\qquad+(n-1){\mathbb{E}}[\log\left\lvert X\right\rvert$].
Essentially the same result holds for discrete random variables, and with the same proof. Ignoring the ${\mathbb{E}}[\log|X|]$ terms and writing $\sigma(X)=H(X_{1}+X_{2})-H(X)$, $\widetilde{\sigma}(X)=H(X_{1}X_{2})-H(X)$, and $\delta(X)=H(X_{1}-X_{2})-H(X)$, we note the following: The discrete analogs of Proposition 7.1 and Theorem 7.2 were proven in [18], Lemma 7.3 holds via the same proof for the discrete case, Lemmas 6.1 and 6.2 are trivial for discrete entropy, and Theorem 6.6 is equivalent to the discrete analog of [12, Theorem 3.11], which holds via the same proof. Therefore, we readily obtain:
Theorem 7.7 (General ring Plünnecke–Ruzsa inequality for discrete entropy).
Suppose the discrete random variables $\{X_{ij}\;;\;1\leq i\leq m,1\leq j\leq n\}$ are i.i.d., taking values in an integral domain, and distributed as $X$. Assuming $X\neq 0$ a.s., we have
$H\left(\sum_{i=1}^{m}\prod_{j=1}^{n}X_{i,j}\right)$ $\leq H(X_{1,1}X_{1,2}\cdots X_{1,n})+(m-1)\big[(n+2)(n-1)\widetilde{\sigma}(X)+(n-1)\delta(X)+\sigma(X)\big]$
$\leq H(X)+\big((m-1)(n+2)+1\big)(n-1)\widetilde{\sigma}(X)+(m-1)(n-1)\delta(X)+(m-1)\sigma(X).$
In the special case when both $\sigma(X)$ and $\widetilde{\sigma}(X)$ are bounded by $\log K$, by the doubling-difference inequality we also have that $\delta(X)\leq 2\log K$, and the discrete version of our general ring Plünnecke–Ruzsa inequality implies the bound
$H\left(\sum_{i=1}^{m}\prod_{j=1}^{n}X_{i,j}\right)\leq H(X)+\big[(n-1)+(m-1)\big(n^{2}+3n-3\big)\big]\log K.$
Interestingly, this is identical to a result independently obtained by Máthé and O’Regan in an updated arXiv preprint version of [18].
7.2 An inequality for slopes
Here we prove an inequality for the entropy for the ratio of sums (or differences) of i.i.d. random variables. It is a partial analog of the discrete entropy bound [18, Theorem 4.7].
Theorem 7.8.
Suppose $X,Y,Z,W$ are i.i.d. continuous random variables, such that $\log\left\lvert X\right\rvert$ and $\log\left\lvert X\pm Y\right\rvert$ are integrable. Then
$h\left(\frac{X\pm Y}{Z\pm W}\right)+5h(X)\leq 4h(XY)+2h(X\pm Y)-3{\mathbb{E}}[\log\left\lvert X\right\rvert]-2{\mathbb{E}}[\log\left\lvert X\pm Y\right\rvert$],
where the inequality holds with either choice of all-plus or all-minus signs, under the corresponding assumption. Equivalently, we have,
$\widetilde{h}\left(\frac{X\pm Y}{Z\pm W}\right)+5\widetilde{h}(X)\leq 4\widetilde{h}(XY)+2\widetilde{h}(X\pm Y).$
Theorem 7.8 immediately follows from the next lemma, upon taking all five random variables to be i.i.d.
Lemma 7.9.
Suppose $X,Y,Z,W$ and $U$ are independent continuous random variables, such that $\log\left\lvert U\right\rvert$, $\log\left\lvert X\pm Y\right\rvert$ and $\log\left\lvert Z\pm W\right\rvert$ are integrable. Then
$h\left(\frac{X\pm Y}{Z\pm W}\right)+h(X,Y,Z,W,U)$ $\leq h(X\pm Y)+h(Z\pm W)+h(UX,UY)+h(UZ,UW)$
$\quad-3{\mathbb{E}}[\log\left\lvert U\right\rvert]-2{\mathbb{E}}[\log\left\lvert Z\pm W\right\rvert$],
where the inequality holds with either choice of all-plus or all-minus signs, under the corresponding pair of assumptions. Equivalently, we have,
$\widetilde{h}\left(\frac{X\pm Y}{Z\pm W}\right)+\widetilde{h}(X,Y,Z,W,U)\leq\widetilde{h}(X\pm Y)+\widetilde{h}(Z\pm W)+\widetilde{h}(UX,UY)+\widetilde{h}(UZ,UW).$
Proof. Let $V_{1}=X\pm Y$ and $V_{2}=Z\pm W$. By Proposition 6.3 applied to $V_{1}$, $1/U$, and $V_{2}$, and Lemma 6.1,
$h(V_{1}/V_{2})+h(U)-2{\mathbb{E}}[\log\left\lvert U\right\rvert]\leq h(V_{1}U)+h(UV_{2})-3{\mathbb{E}}[\log\left\lvert U\right\rvert]-2{\mathbb{E}}[\log\left\lvert V_{2}\right\rvert$].
Rearranging and applying Proposition 7.1 to $h(V_{1}U)$ and $h(UV_{2})$, we find
$h\left(\frac{X\pm Y}{Z\pm W}\right)+h(U)$ $\leq h(U,X\pm Y)+h(U,Z\pm W)+h(UX,UY)+h(UZ,UW)$
$\quad-h(U,X,Y)-h(U,Z,W)-3{\mathbb{E}}[\log\left\lvert U\right\rvert]-2{\mathbb{E}}[\log\left\lvert Z\pm W\right\rvert$].
Using the independence assumption and rearranging yields the desired inequality.   $\Box$
8 On the entropic sum-product phenomenon in $\mathbb{Z}$}
For a subset $A$ of the integers $\mathbb{Z}$}, let $A\cdot A$ denote the product set $A\cdot A=\{a_{1}a_{2}:a_{1},a_{2}\in A\}$. In 1983, Erdős and Szemerédi [3] observed that $A+A$ and $A\cdot A$ cannot both be small simultaneously. Specifically, they showed that there is a positive constant $\varepsilon$ such that, for $A\subset{\mathbb{Z}}$,
$\max\left\{\left\lvert A+A\right\rvert,\left\lvert A\cdot A\right\rvert\right\}\geq\left\lvert A\right\rvert^{1+\varepsilon-o(1)},$ (58)
where the $o(1)$ term tends to zero as $|A|\to\infty$. Although they did not provide an explicit estimate for $\varepsilon$, it is widely conjectured – though still unproven – that the result should hold for any $\varepsilon\leq 1$. The best result to date is by Bloom [2], with $\varepsilon=\frac{1}{3}+\frac{2}{951}$; this result also holds more generally for finite $A\subset{\mathbb{R}}$.
In view of the entropy/cardinality correspondence discussed in the Introduction, it might be tempting to conjecture an analogous result for the entropy. Namely, that for i.i.d. $\mathbb{Z}$}-valued random variables $X,X^{\prime}$,
$\max\{H(X+X^{\prime}),H(X\cdot X^{\prime})\}\geq(2-o(1))H(X),$
as $H(X)\to\infty$, or at least that, as conjectured by Goh [5],
$\max\{H(X+X^{\prime}),H(X\cdot X^{\prime})\}\geq(1+\epsilon-o(1))H(X),$ (59)
for some $\varepsilon>0$. In this section we show that, if such a result were to hold, it would necessarily be with $\varepsilon\leq 1/3$. In particular, the obvious analog of Bloom’s bound fails for the entropy.
Example 8.1.
For a large positive integer $n$ that we will send to infinity, consider the discrete real-valued random variable $X$ which takes value 0 with probability $\frac{1}{3}$, and with probability $\frac{2}{3}$ takes a value uniformly at random in $\{1,2,\dots,n\}$. Asymptotically as $n\to\infty$, its entropy is $H(X)=\frac{2}{3}\log n+O(1)$. Let $X^{\prime}$ be an independent copy of $X$, and let $A$ and $A^{\prime}$ be the indicators for $X=0$ and $X^{\prime}=0$, respectively. Note that $H(X+X^{\prime}|A=0,A^{\prime}=0)=\log n+O(1)$, because
$H(X+X^{\prime}|A=0,A^{\prime}=0)\geq H(X+X^{\prime}|A=0,A^{\prime}=0,X^{\prime})=H(X|A=0)=\log n$
and
$H(X+X^{\prime}|A=0,A^{\prime}=0)\leq\log\left\lvert\left\{1,2,\dots,n\right\}+\left\{1,2,\dots,n\right\}\right\rvert=\log(2n-1)\leq\log n+\log 2.$
Thus,
$H(X+X^{\prime})=H(X+X^{\prime}|A,A^{\prime})+O(1)=\left(1-\frac{1}{9}\right)\log n+O(1)=\frac{8}{9}\log n+O(1),$
where the first equality follows from the inequalities
$H(X+X^{\prime}|A,A^{\prime})\leq H(X+X^{\prime})$ $\leq H(X+X^{\prime},A,A^{\prime})$
$=H(X+X^{\prime}|A,A^{\prime})+H(A,A^{\prime})$
$\leq H(X+X^{\prime}|A,A^{\prime})+\log 4.$
We also observe
$H(X\cdot X^{\prime}|A=0,A^{\prime}=0)\leq\log\left\lvert\left\{1,2,\dots,n\right\}\cdot\left\{1,2,\dots,n\right\}\right\rvert\leq\log(n^{2}),$
and so
$H(X\cdot X^{\prime})=H(X\cdot X^{\prime}|A,A^{\prime})+O(1)\leq\frac{4}{9}\log(n^{2})+O(1)=\frac{8}{9}\log n+O(1),$
so
$\max\left\{H(X+X^{\prime}),H(X\cdot X^{\prime})\right\}\leq\left(\frac{4}{3}+o(1)\right)H(X).$
By taking $n\to\infty$, we can obtain arbitrarily large $H(X)$, so this example shows (59) cannot hold in general for any $\varepsilon>\frac{1}{3}$.
Note that the choice of the value 1/3 for ${\mathbb{P}}(X=0)$ is optimal in that any other value gives a constant larger than $4/3$ in the final result. Therefore, this method cannot improve our estimate that, if (59) were to hold, it could only hold for some $\varepsilon\leq 1/3$.
However, we know (58) holds for some $\varepsilon>\frac{1}{3}$, and one may wonder whether we can use this inequality to show (59) at least in the case when $X$ is uniformly distributed over some finite set $A\subset{\mathbb{R}}$. To that end, one may wonder whether there are $\varepsilon,\varepsilon^{\prime}\in(0,1)$ such that, for any finite set $A\subset{\mathbb{R}}$,
$\left\lvert A+A\right\rvert\geq\left\lvert A\right\rvert^{1+\varepsilon}\quad\implies\quad H(U_{A}+U_{A}^{\prime})\geq(1+\varepsilon^{\prime}-o(1))H(U_{A}),$
where $U_{A}$ and $U_{A}^{\prime}$ are i.i.d. random variables uniformly distributed on $A$. The weakest version of this result would be of the form, $\left\lvert A+A\right\rvert\geq\left\lvert A\right\rvert^{2-\varepsilon}$ implies $H(U_{A}+U_{A}^{\prime})\geq(1+\varepsilon-o(1))H(U_{A})$, for some small $\varepsilon>0$”. The following example, a slight generalization of one in [25], shows that this does not hold for any $\varepsilon>0$.
Example 8.2.
Let $n$ be a large positive integer and $\varepsilon>0$. Take $A$ be the union of $\left\{1,2,\dots,n\right\}$ and a set $B$ consisting of $n^{1-\varepsilon/2}$ other integers in general position, i.e., such that there are no nontrivial pairwise sum relations among themselves and with $\left\{1,2,\dots,n\right\}$: If $a+b=c+d$ for $a,c,d\in A$ and $b\in B$, then $\{a,b\}=\{c,d\}$.
For large $n$ we have $\left\lvert A\right\rvert\sim n$, while,
$\left\lvert A+A\right\rvert\geq n^{2-\varepsilon/2}\geq\left\lvert A\right\rvert^{2-\varepsilon},$
because of the sums between $\left\{1,2,\dots,n\right\}$ and the $n^{1-\varepsilon/2}$ integers in $B$. Let $U_{A}$ and $U_{A}^{\prime}$ be independent and uniformly distributed on $A$, and write $Z$ and $Z^{\prime}$ for the indicators of the events, $\{U_{A}\in\left\{1,2,\dots,n\right\}\}$ and $\{U_{A}^{\prime}\in\left\{1,2,\dots,n\right\}\}$, respectively. Note that $|A+A|\leq|A|^{2}$ and ${\mathbb{P}}(Z=1),{\mathbb{P}}(Z^{\prime}=1)\to 1$ as $n\to\infty$. Then, $H(U_{A}+U_{A}^{\prime})\sim H(U_{A}+U_{A}^{\prime}|Z,Z^{\prime})$, where
$H(U_{A}+U_{A}^{\prime}|Z,Z^{\prime})\leq{\mathbb{P}}(Z=1,Z^{\prime}=1)(\log n+O(1))+(1-{\mathbb{P}}(Z=1,Z^{\prime}=1))\log\left\lvert A+A\right\rvert\sim\log n.$
Therefore, despite having $\left\lvert A+A\right\rvert\geq\left\lvert A\right\rvert^{2-\varepsilon}$, we still have $H(U_{A}+U_{A}^{\prime})<(1+\varepsilon-o(1))H(U_{A})$.